OpenStudy (anonymous):

Define f(x)= x-4, 1≤x<2 1/x-3, 2≤x<5 -x+5.5, 5≤x a) show that f(x) is continuous at x=5 b) where on the interval [1,4] is f discontinuous?

OpenStudy (anonymous):

\[f(x)=\begin{cases}x-4&\text{for }1\le x<2\\\\\\\frac{1}{x-3}&\text{for }2\le x<5\\\\\\ \frac{11}{2}-x&\text{for }5\le x\end{cases}\] Is this the function?

OpenStudy (anonymous):

everything is the same except for the last one it is -x+5.5

OpenStudy (anonymous):

Right, well 5.5 is 11/2.

OpenStudy (anonymous):

oh yea lol

OpenStudy (anonymous):

Anyway, to show \(f(x)\) is continuous at \(x=5\), you have to show that \[\lim_{x\to5^-}f(x)=\lim_{x\to5^+}f(x)\]

OpenStudy (anonymous):

oh then how do I show that?

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles And that \(f(5)\) exists and is equal to these limits, but that's given by the function. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

\[\lim_{x\to5^-}\frac{1}{x-3}=\lim_{x\to5^+}\left(\frac{11}{2}-x\right)\] Is this true?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Okay, so \(f\) is indeed continuous at 5.

OpenStudy (anonymous):

For the second part, think about what values in the desired interval make the function undefined.

OpenStudy (anonymous):

but what does it mean by interval [1,4]?